THE PERIMETER OF AN EQUILATERAL TRIANGLE IS 12CM


A= WHAT IS THE LENGTH OF EACH SIDE OF THE TRIANGLE


B= WHAT IS THE SIZE OF EACH ANGLE OF THE TRIANGLE


1
Expert's answer
2021-09-13T00:04:46-0400

A) "P=3a\to a=\frac{P}{3}=\frac{12}{3}=4 cm."

B) "\alpha=\frac{180^0}{3}=60^0."